Please help me answer this word problem:
When you add three to a number and then divide the sum by seven, you get the same result as when you add three to the same number and then divide the sum by twelve. What is the number?

Thank you for your time!

Answer by rfer(16322)

(x+3)/7=(x+3)/12
12x+36=7x+21
12x-7x=21-36
5x=-15
x=-15/5
x=-3
-------
(-3+3)/7=(-3+3)/12
0=0